### v4

# Nested homotopy models of finite metric spaces and their spectral homology

Abstract.
For a real we consider the notion of -homotopy equivalence in the category quasimetric spaces, which includes metric spaces and directed graphs. We show that for a finite quasimetric space there is a unique (up to isometry) -homotopy equivalent quasimetric space of the minimal possible cardinality. It is called the -minimal model of . We use this to construct a decomposition of the magnitude-path spectral sequence of a digraph into a direct sum of spectral sequences with certain properties.
We also construct an -homotopy invariant of a quasimetric space called spectral homology, that generalizes many other invariants: the pages of the magnitude-path spectral sequence, including path homology, magnitude homology, blurred magnitude homology and reachability homology.
1. Introduction
1.1. Background
For over a decade, two theories have been actively developed. The first theory is the theory of magnitude of metric spaces and their magnitude homology. This theory was developed by Leinster, Meckes, Shulman, Hepworth, Willerton, Yoshinaga, Asao and others [leinster2013magnitude, meckes2013positive, leinster2019magnitude, hepworth2015categorifying, leinster2021magnitude, kaneta2021magnitude, asao2021geometric]. It explores an invariant of a metric space called magnitude. On one hand, the magnitude is a real number, which is a continuous analogue of the cardinality of a finite set. On the other hand, it exhibits some properties similar to those of the Euler characteristic. The concept of magnitude is also related to the notion of diversity [leinster2016maximizing], [leinster2021entropy]. Subsequently it was discovered that it is possible to introduce a homology theory for metric spaces, so that for finite metric spaces the Betti numbers of this theory are related to the magnitude in the similar way as ordinary Betti numbers are related to the Euler characteristic. This homology theory is called magnitude homology.
The second theory is the GLMY-theory of digraphs. This theory was developed by Grigorβyan, Lin, Muranov, Yau and others [grigor2012homologies, grigor2014homotopy, grigor2017homologies, grigor2020path, grigor2014graphs, grigor2018path, grigor2018path2, grigor2022advances, ivanov2022simplicial]. It implements a homotopical approach to the theory of digraphs. Its central concept is the path homology of a digraph, also known as GLMY-homology. The path homology behaves in many ways similar to the homology of spaces. For example, there is an analogue of the KΓΌnneth formula; there is a notion of suspension of a digraph that shifts the path homology; there is a notion of the fundamental group of a digraph, whose abelianization equals the first path homology group, and so forth. The theory also has the notion of homotopy of morphisms in the category of digraphs, and path homology is homotopy invariant. Moreover, there is a theory of covering digraphs, analogous to the theory of covering spaces in algebraic topology [di2023path].
Asao later showed that for the case of digraphs, these two theories could be unified into a single theory [asao2023magnitude],[asao2023magnitude2]. He showed that each digraph defines a spectral sequence , whose first page is the magnitude homology and the second page contains the path homology as the zeroth row. This spectral sequence is called magnitude-path spectral sequence. He also defines the concept of -homotopy on the category of digraphs and shows that is an -homotopy invariant. For , this concept reduces to homotopy as defined in the GLMY-theory. Thus, this result generalizes the fact that the path homology is a homotopy invariant.
For a digraph , one can define the distance between vertices as the length of the shortest path from to . If there is no path from to , we assume Note that this distance is not necessarily symmetric and can be equal to infinity. However, it satisfies the triangle inequality and if and only if Such generalizations of metric spaces are referred to as quasimetric spaces. We will work in the category of quasimetric spaces where morphisms are short maps.
The aim of our work is to study the concept of -homotopy, introduced by Asao, for the case of quasimetric spaces and any real We study this concept from two perspectives. First, we study finite quasimetric spaces that have the least cardinality among quasimetric spaces in a given -homotopy equivalence class. Secondly, we introduce a new invariant of the -homotopy equivalence for any quasimetric spaces, and study this invariant. Both parts are related to the magnitude-path spectral sequence: in the first part we obtain its decomposition into a direct sum of some smaller spectral sequences; in the second part we generalize its pages to the case of any quasimetric spaces.
1.2. Nested -minimal models
In this theory, after the -homotopy equivalence of short maps is defined, -homotopy equivalences and -deformation retracts are defined similarly to how it is defined in the classical algebraic topology. We say that a finite quasimetric space is -minimal, if it has no proper -deformation retracts. A quasimetric space is called the -minimal model of a quasimetric space if is -minimal and it is -homotopy equivalent to We prove the following theorem.
Theorem.
Let be a real number. Then the following holds.
Any finite quasimetric space has an -minimal model, which is unique up to isometry.
Two finite quasimetric spaces are -homotopy equivalent if and only if their -minimal models are isometric.
Any finite quasimetric space has an -deformation retract, which is its -minimal model.
The -minimal model of a space is denoted by Note that we do not claim that is a functor, and the inclusion and the retraction are not uniquely defined. It is easy to see that if then Using this we obtain that there is a finite number of real numbers such that for We call these points homotopy jumping points and set
Thus for any finite quasi-metric space we obtain a nested sequence of retracts
Unfortunately, the map from the set of finite subsets of to the set of finite subsets of is not continuous with respect to the Hausdorff distance (Example 2.12).
If is a digraph, the sequence (1.2) is a sequence of convex subdigraphs, such that is a retract of Using this we obtain a decomposition of the magnitude-path spectral sequence into a direct sum of spectral sequences
such that and for any we have
However, we do not claim that this decomposition is functorial by .
We say that a finite digraph is homotopy stable, if for any Any finite digraph defines a homotopy stable digraph by the formula where is greater then all homotopy jumping points of . The decomposition (1.3) shows that the magnitude-path spectral sequence contains a summand corresponding to the homotopy stable digraph such that for all big enough This indicates that homotopy stable digraphs may be an important subject for further study.
1.3. Spectral homology
The magnitude-path spectral sequence is defined for the discrete case of digraphs and its first page is the magnitude homology. However the magnitude homology is defined for all quasimetric spaces. Spectral homology is an attempt to define the magnitude-path spectral sequence for any quasimetric space so that its βfirst pageβ is the magnitude homology. In some sense the spectral homology is a βcontinuousβ generalization of the magnitude-path spectral sequence. In order to define the spectral homology of a quasimetric space, we need to define the spectral homology of an -filtered chain complex so that the spectral homology is a generalization of pages of the spectral sequence associated with a -filtered chain complex. In other words, the spectral homology of an -filtered complex is an analogue of the spectral sequence of a -filtered chain complex.
For an -filtered chain complex over a ring a real number a non-empty interval and a natural number we define a -module called spectral homology, or -spectral homology, if we want to specify The spectral homology generalizes the spectral sequence associated with a -filtered chain complex in the following sense. If is a -filtered chain complex, then its -filtration can be extended to an -filtration by the formula for In this case for a natural the pages of the associated spectral sequence can be presented as the spectral homology
Let us give the definition of A left infinite interval in is an interval of the form For a left infinite interval we set If we present a non-empty interval as the difference of left infinite intervals then
Moreover, there is another more technical equivalent definition of the spectral homology that shows that for the module is a subquotient of (see details in Subsection 4.2).
The magnitude-path spectral sequence of a digraph is defined using a -filtered chain complex which is called reachability chain complex (see [hepworth2023reachability]). This complex is defined similarly for the case of quasimetric space However, in this case we obtain an -filtration on the chain complex Then the spectral homology of is defined as the spectral homology of the -filtered chain complex
We show that is an -homotopy invariant.
The spectral homology is natural by the interval in some sense. In particular, there is a homomorphism for This allows us to define a persistent module such that
We prove that some other types of homology can be presented as spectral homology.
Theorem.
Let be a quasimetric space and be a digraph. Then the following holds.
The magnitude homology can be presented as -spectral homology
The persistent module of blurred magnitude homology (see [govc2021persistent]) can be presented as the persistent -spectral homology
The reachability homology (see [hepworth2023reachability]) can be presented as -spectral homology for any
For any integer the -th page of the magnitude-path spectral sequence of can be presented as -spectral homology
In particular, the path homology can be described as
In the end of the paper we give a description of some low-dimensional spectral homology. In particular, for we give a description of in terms of adjacent pairs, similar to the description of given in [leinster2021magnitude, Cor. 4.5]. We also work out the following continuous example. Let be an arc in the circle of radius (with the Euclidean distance) that does not lie in one semicircle.
If we assume that and then
So the first spectral homology detects the distance for the βsingular pairβ and shows that the arc is not -contractible for Moreover, in this case has finite rank, while has infinite rank.
1.4. Acknowledgments
The author is grateful to Lev Mukoseev for fruitful discussions and Example 2.8.
Contents
- 1 Introduction
- 2 -minimal models of quasimetric spaces
- 3 Magnitude-path spectral sequence of a digraph
4 Spectral homology
- 4.1 Reminder on spectral sequence of a -filtered chain complex
- 4.2 Spectral homology of an -filtered chain complex
- 4.3 -homotopy of -filtered chain complexes
- 4.4 Spectral homology of a quasimetric space
- 4.5 Relation of spectral homology with other types of homology
- 4.6 Low-dimensional spectral homology
2. -minimal models of quasimetric spaces
2.1. Quasimetric spaces
A quasimetric space is a set equipped by a map such that the triangle inequality holds for any and if and only if Note that we allow the distance to be equal to infinity, and it is not necessary symmetric, but the distance between distinct points is positive. Two sources of examples of quasimetric spaces are metric spaces and directed graphs with the shortest path distance.
A map between two quasimetric spaces is called short (or 1-Lipschitz, or distance-decreasing), if
for any The set of all short maps from to is denoted by The category of quasimetric spaces with short maps as morphisms is denoted by Note that isomorphisms in this category are isometries (distance preserving bijections).
For two short maps we set
It is easy to see that this defines a structure of quasimetric space on Note that for any short maps and we have
2.2. -homotopy category
Let be a real number. We say that two points of a quasimetric space are -homotopic, denoted by if there is a sequence of points of such that for each we have either or In other words, is the minimal equivalence relation such that implies We say that two short maps are -homotopic if they are -homotopic as points in
Lemma 2.1.
If and are short maps between quasimetric spaces and then In other words, is a congruence relation on the category
Proof.
Follows from (2.3). β
Since is a congruence relation on we can define the -homotopy category of quasimetric spaces as the quotient-category . We say that a short map is an -homotopy equivalence, if its image in is an isomorphism, and two quasimetric spaces are -homotopy equivalent, if they are isomorphic in
For a quasimetric space we denote by the monoid of endomorphisms of in We will also denote by the submonoid consisting of short maps such that
2.3. -deformation retracts
If is a quasimetric space and is a subset, will be treated as a quasimetric space with the induced metric. We will denote by the canonical embedding. A retract of a quasimetric space is a subset such that there exists a short map , called retraction, such that . We say that the retract is proper, if An -deformation retract is a retract with a retraction such that
Proposition 2.2.
Let be a finite quasimetric space and . Then there exists a natural such that is idempotent, and for any such the subset is an -deformation retract of
Proof.
Since is finite, there exists such that is an idempotent. Lemma 2.1 implies that Therefore Set Then we can present as the composition where is the canonical embedding and is some surjection. Since the quasimetric on is induced from we obtain that is a short map. Note that Since is idempotent, we have Using that is injective, and is surjective, we obtain β
2.4. -minimal spaces
We say that a finite quasimetric space is -minimal, if it has no proper -deformation retracts.
Proposition 2.3.
Let be a finite quasimetric space. Then the following statements are equivalent.
is -minimal;
any such that either or is an automorphism.
Proof.
Take Proposition 2.2 implies that there exists such that is an idempotent and is an -deformation retract. Since is -minimal, we obtain Using that is finite, we obtain that is a bijection and an isometry. Combining this with the fact that an idempotent, we obtain Therefore is an automorphism.
Obvious.
Consider an -deformation retract with a retraction We need to prove that Since there exists a sequence such that for any either or Prove by induction that is an automorphism. The base of induction is obvious. Assume that is an automorphism. If then and hence is an automorphism. If then and hence, is an automorphism. It follows that is an automorphism. Since is an automorphism, β
Proposition 2.4.
Let and be -minimal finite quasimetric spaces. If is an -homotopy equivalence, then is an isometry.
Proof.
Since is an -homotopy equivalence, there is a short map such that and By Proposition 2.3 we obtain that and are automorphisms. Therefore is an isomorphism. β
2.5. -minimal model of a space
We say that is an -minimal model of a finite quasimetric space if is -minimal, and is -homotopy equivalent to
Theorem 2.5.
Let be a real number. Then the following holds.
Any finite quasimetric space has an -minimal model, which is unique up to isometry.
Two finite quasimetric spaces are -homotopy equivalent if and only if their -minimal models are isometric.
Any finite quasimetric space has an -deformation retract, which is its -minimal model.
Proof.
For a finite quasimetric space we take any -deformation retract of the minimal possible cardinality. Denote the retraction by We claim that is -minimal. Indeed, if is an -deformation retract with a retraction then the equations and imply and relations and imply Hence is an -deformation retract of which is subset of and using the minimality of the cardinality of we obtain
Proposition 2.4 implies that the -minimal model is defined uniquely up to isometry. It also implies that -homotopy equivalent quasimetric spaces have isomorphic -minimal homotopy models. β
The -minimal model of a quasimetric space is denoted by Note that for we have
Example 2.6.
Let be a finite metric space (with symmetric finite distance). For the metric space is a one-point space. Here we see that the embedding is not uniquely defined, because we can take any map from the one-point space to as the inclusion.
Example 2.7.
Consider the following graph
as a metric space with the shortest path distance. Take Then the -minimal model is the pentagon with the vertex set There are two choices of retractions for this -deformation retract and In particular, this example shows that the retraction canβt be defined by a universal property in .
Example 2.8.
The author is grateful to Lev Mukoseev for the following non-trivial but clarifying example. Consider the following digraph with four vertices together with a subdigraph with three vertices
It is easy to see that is a retract of because there is a retraction such that We claim that is a -deformation retract of and In particular
In order to prove it consider an endomorphism defined by
It can be depicted as follows.
Then and It follows that
Remark 2.9.
We say that a finite metric space is in general position if for any two pairs of distinct points and such that the distances are not equal Then for any finite metric space in general position, if consists of more then one point, then the embedding is uniquely defined.
2.6. Homotopy jumping points
For a finite quasimetric space we consider a function from non-negative real numbers to the interval defined as the ratio of cardinalities
Lemma 2.10.
The function is decreasing, lower semi-continuous and it has a finite number of values.
Proof.
If then and hence Now assume that and prove that Since is decreasing, we can assume that is an decreasing sequence. Since is finite, for we have that for any two points either or Therefore, for the inequality is equivalent to the inequality Hence for we have and β
Lemma 2.10 implies that has a finite number of jumping points such that for any We will call them homotopy jumping points of and set
If we also set and then
Remark 2.11.
Let us call a quasimetric space homotopy stable, if In other words, is homotopy stable, if for any If we set then Proposition 2.3 says that is homotopy stable if and only if For example, the quasimetric spaces defined by the digraphs
(2.14) | Β Β Β Β Β Β Β Β Β |
are homotopy stable.
Any finite quasimetric space with the set of homotopy jumping points defines a homotopy stable finite quasimetric space given by
if and if For a finite quasimetric space with a point such that for any the metric space is always a one-point space.
Example 2.12 (Discontinuity of with respect to the Hausdorff distance).
For any we consider a subset of of six points
which can be depicted as follows.
Note that with respect to the Hausdorff distance. For we have because there is a retraction defined by such that It is also easy to see that is a one-point set. Using this we get
However, for the direct verification using brute force shows that for any non-identical short map we have A non-identical short map with the minimal possible is the map that sends to and sends to Using this we obtain
Since , the map is not continuous with respect to the Hausdorff distance.
3. Magnitude-path spectral sequence of a digraph
3.1. -minimal models of digraphs
We define a digraph as a pair where such that for any Elements of are called vertices, elements of are called arrows, arrows of the form are called degenerate arrows. A morphism of digraphs is a map such that implies The category of digraphs is denoted by A path in a digraph is a tuple such that Note that any morphism sends a path to a path.
Any digraph defines a quasimetric space, that we denote by whose set of points is and the distance from to is defined as the infimum of the lengths of paths from to Therefore we get a functor
It is easy to check that the functor is full and faithful. So we can identify with a full subcategory of So we can define -homotopic maps for digrphs.
A subdigraph of a digraph is called convex, if the shortest path distance on coincides with the induced distance: for any vertices of A subdigraph is called a retract, if there is a morphism of digraphs such that where is the canonical embedding.
Lemma 3.1.
A retract of a digraph is a convex subdigraph.
Proof.
Let be a retract of It is obvious that for any On the other hand, for any path from to in we have a path from to in Hence β
Lemma 3.2.
Let be a digraph and be a retract of Then for a uniquely defined retract
Proof.
Consider a subdigraph such that and Denote by a retraction. If then and hence So defines a morphism of digraphs Therefore is a retract of In particular is a convex subdigraph. It follows that the distance between points in coincides with the distance in β
Lemma 3.2 implies that for a finite digraph and a natural there is a digraph such that that can be included into as an -deformation retract, but the embedding is not unique. Therefore for a finite digraph we obtain a nested sequence of convex sudigraphs
where the inclusions are not uniquely defined, but the digraphs are defined uniquely up to isomorphism.
3.2. Reminder on the magnitude-path spectral sequence
Here we remind the construction and the main properties of the magnitude-path spectral sequence of a digraph that can be found in [asao2023magnitude], [asao2023magnitude2], [hepworth2023reachability],[di2023path].
Let be a commutative ring. For a set we will denote by the free -module freely generated by For a digraph we consider a filtered chain complex over called reachability complex, whose -th component is freely generated by tuples of vertices without consecutive repetitions such that
Tuples with a consecutive repetition will be identified with zero in the chain complex. Then the differential is defined by the formula
If we set the filtration on the chain complex is defined so that for an integer the component is generated by tuples such that
The spectral sequence associated with this filtered chain complex is called magnitude-path spectral sequence of and will be denoted by Its first page consists of the magnitude homology
the zero row of the second page is the path homology
Since the filtration of is exhaustive and bounded below, the spectral sequence converges to the homology of the chain complex [weibel1995introduction, Th.5.5.1(2)] (here by convergence we mean βunboundedβ convergence in the sense of unbounded spectral sequences [weibel1995introduction, Β§5.2.11])
The homology of this chain complex is known as reachability homology and denoted by The reachability homology is equal to the homology of the reachability preorder which is defined so that its elements are points of and if and only if
The -st page of the magnitude-path spectral sequence is an -homotopy invariant in the following sense. If two -homotopy invariant morphisms, then they induce equal morphisms on the -st pages
(see [asao2023magnitude2, Th. 4.37]).
3.3. Decomposition of the magnitude-path spectral sequence
Theorem 3.3.
Let be a finite digraph and let be the set of homotopy jumping points of . Then the magnitude-path spectral sequence can be decomposed into a direct sum of spectral sequences
such that and for any we have
In particular,
Proof.
It is easy to see that, if is a spectral sequence and is an idempotent morphism of spectral sequences, then can be decomposed into a sum of spectral sequences In other words, the category of spectral sequences is Karoubian. Therefore, if is a retract of a spectral sequence then is a direct sum of spectral sequences
If then there is nothing to prove. Assume Chose a sequence of nested minimal models
with retractions for Since is a retract in we obtain that is a retract of Hence Then we obtain a decomposition (3.11) with the property (3.12). Since the embedding is an -homotopy equivalence, we obtain that Therefore β
Remark 3.4.
We do not claim that the decomposition in Theorem 3.3 is natural.
4. Spectral homology
4.1. Reminder on spectral sequence of a -filtered chain complex
In this subsection we remind an explicit construction of pages of the spectral sequence associated with a filtered chain complex that can be found in [cartan1999homological, Ch.XV,Β§1]. We will use the following lemma from the book of Cartan-Eilenberg.
Lemma 4.1 ([cartan1999homological, Ch.XV,Lemma 1.1]).
For any commutative diagram of -modules of the following form, in which the row is exact
there is an inclusion and the map induces an isomorphism
Let be a commutative ring and be a -filtered chain complex over with an increasing filtration
For the sake of simplicity we set
Consider the associated spectral sequence Here we will not discuss the convergence of the spectral sequence, so we do not need any assumptions on the filtration. The first page of this spectral sequence is described as
Let us describe all pages as subquotients of the first page. For any natural we consider the following short exact sequences induced by the embeddings
and
and the associated long exact sequences of homology. Further we set
and
Then the -th page of the spectral sequence can be described as
Using the morphism of short exact sequences,
we obtain a commutative diagram, in which the row is exact.
Applying Lemma 4.1 to this diagram, we obtain another description of the (see [cartan1999homological, Ch.XV,Β§1,(8)])
4.2. Spectral homology of an -filtered chain complex
A subset of the set of real numbers is called an interval if it is convex. We say that a interval is left infinite, if for any and we have There are four types of left infinite intervals: and Any non-empty interval can be uniquely presented as the difference of left infinite intervals where
For example and We define an order on the set of non-empty intervals such that if and only if and For example,
For an interval and a non-negative real number we also set and
Note that
Assume that is an -filtered chain complex over a ring . So for any we have a chain subcomplex such that implies When is fixed we will omit it in notations For a left infinite interval we set
Further, for a non-empty interval we set
If we have a morphism
induced by inclusions and . Moreover, if this morphism is a monomorphism, and, if it is an epimorphism. If we have three distinct left infinite intervals then these morphisms define a short exact sequence of chain complexes
The -spectral homology of is defined as follows
Further for any we will define -spectral homology as a subquotient of
For any we consider short exact sequences of chain complexes of the form (4.20) similar to (4.6), (4.7)
and the associated long exact sequences of homology. Then we define -spectral cycles and -spectral boundaries as
and
Proposition 4.2.
There is an inclusion
Moreover, if we set
then the map induces an isomorphism
Proof.
The -module defined in Proposition 4.2 is called -spectral homology of the -filtered chain complex A morphism of -filtered chain complexes is a chain map such that It is easy to see that the construction is natural and defines a functor
Corollary 4.3.
There are isomorphisms
Proposition 4.4.
Let be a -filtered chain complex. We extend the -filtration to an -filtration on by the formula
Then for a natural number we have
Proof.
For we obtain and Therefore, Then the short exact sequences (4.6), (4.7) coincide with short exact sequences (4.22), (4.23).
Similarly for We obtain and Therefore β
Proposition 4.5.
Let be an -filtered complex such that Then for any we have
Proof.
In this case and Therefore and The assertion follows. β
Remark 4.6 (Naturalness in the interval and persistent spectral homology).
Note that, if then and Therefore, the commutative diagram
defines a morphism
Moreover, it is easy to see that for we have In particular, it follows that we can consider persistent module such that
Proposition 4.7.
For any there is an isomorphism
Moreover,
and hence, is a subquotient of
Proof.
If then and Then and are images of the horizontal arrows in the following commutative diagram.
Therefore the image of is equal to the image of the diagonal map
The inclusion follows from the fact that the map factors as The second inclusion can be proved similarly. β
4.3. -homotopy of -filtered chain complexes
Similarly to [cartan1999homological, Ch.XV,Β§3] for we say that two morphisms of -filtered chain complexes are -homotopic, if there is a morphism of graded modules of degree one such that and
Proposition 4.8 (cf. [cartan1999homological, Ch.XV,Prop.3.1]).
Any -homotopic morphisms of -filtered chain complexes induce equal morphisms on the -spectral homology
Proof.
We present as the difference of left infinite intervals We set and Then
Here we denote by and the morphisms of chain complexes induced by embeddings
Set and consider the commutative diagram
Then it is sufficient to prove that Take an element that represents an element from Then Since we have
Using that is an -homotopy and we obtain Hence represents a zero element in Using that is an -homotopy once again, we obtain Therefore represents a zero element in too. It follows that represents zero element in So we obtain β
4.4. Spectral homology of a quasimetric space
Let be a commutative ring and be a quasimetric space. Define the -filtered reachability chain complex similarly to the case of digraphs. Its -th component is freely generated by tuples without consecutive repetitions such that
In this complex we identify sequences with consecutive repetitions with zero. Then the differential is defined by the formula
If we set the -filtration is defined so that for a real number the component is generated by tuples such that
For a non-empty interval a real number and natural the spectral homology of is defined as the spectral homology of this -filtered chain complex
It is easy to see that these constructions are natural and they define functors
By Remark 4.6 the spectral homology is natural by the interval. In particular, we can consider the persistent spectral homology which is a persistent module such that
Theorem 4.9.
Any -homotopic morphisms of quasimetric spaces induce equal morphisms on -spectral homology
for any and .
Proof.
The differential in the reachability complex is defined by the formula where
Set and For we consider the map given by
These maps satisfy simplicial homotopy relations
Therefore, if we set we obtain
If then
It follows that is an -homotopy between morphisms -filtered chain complexes. Hence the statement follows from Proposition 4.8. β
Corollary 4.10.
If and are -homotopy equivalent quasimetric spaces, then
In particular, if is finite, then
4.5. Relation of spectral homology with other types of homology
Let be a quasimetric space. Then in our notation its magnitude homology can be defined as
and the blurred magnitude homology [govc2021persistent] are defined as a persistent module such that
Theorem 4.11.
Let be a quasimetric space and be a digraph. Then the following holds.
The magnitude homology can be presented as -spectral homology
In particular, for any the group is an -homotopy invariant subquotient of
The persistent module of blurred magnitude homology (see [govc2021persistent]) can be presented as persistent -spectral homology
In particular, for any the persistent module is an -homotopy invariant subquotient persistent module of
The reachability homology (see [hepworth2023reachability]) can be presented as -spectral homology
for any
For any integer the -th page of the magnitude-path spectral sequence of can be presented as -spectral homology
In particular, path homology can be described as
4.6. Low-dimensional spectral homology
In this section we fix a quasimetric space and we will omit it in some notations. For an interval we denote by the equivalence relation on such that if and only if there is such that
Proposition 4.12.
Let be an interval and be a real number. Then, assuming we have
In particular,
If then
Proof.
We set If then and If we have and Therefore is spanned by differences where It follows that Hence β
A pair of points of a quasimetric space is called -adjacent, if and there is no a point such that and The set of -adjacent pairs is denoted by Then it is known [leinster2021magnitude, Cor. 4.5] that
We will prove a generalization of this statement describing for
Let Denote by the set of all pairs of points such that
We consider the minimal equivalence relation on such that the following holds
β’
if and
β’
if and
We say that a pair is -trivial, if there exists a point such that
In particular, if then is -trivial. We denote the set of -trivial pairs by
We say that a pair is homotopy -trivial, if it is equivalent to a trivial pair with respect to the equivalence relation defined above. The set of homotopy -trivial pairs is denoted by
A pair is called -adjacent, if it is not homotopy -trivial
We consider the restricted equivalence relation on
Proposition 4.13.
Let be a quasimetric space and Then
Proof.
Set Then
Consider the map defined so that it sends pairs from to zero, and all other pairs to their equivalence classes It is easy to see that is surjective. We need to prove that
Let us prove that If then either or because if and then Assume first that Then we have four cases.
Case 1.1: and Then
Case 1.2: and Then In this case and hence
Case 1.3: and Then and In this case and In particular, and
Case 1.4: and Then and In this case and
If we have symmetric cases. So we proved that
Let us prove that First we need to check that if and then Here we assume that a pair such that represents the zero element in and a triple such that represents the zero element in
Let such that and Then and Therefore The case of pairs such that and is similar. It follows that the differences of equivalent pairs from are in the image of
Now assume that Then there exists such that and Therefore and It follows that β
Theorem 4.14.
Let be a quasimetric space and Then there is an isomorphism
where is the equivalence relation induced from the equivalence relation on
Proof.
We set Since we have Then we obtain that any element in is a cycle. It follows that
Therefore, is a submodule of generated by pairs such that Then the statement follows from Proposition 4.13. β
Assume that Then we consider the following filled ellipsoid of revolution with foci
We will also consider open balls of radius
We define the -thick open interval between as
It can be depicted as the following gray area between and .
So for such that we have if and only if
Corollary 4.15.
Let and Assume that for each such that the intersection of with is not empty. Then
Proof.
In this case all elements from are in and β
Assume that We say that a (continuous) path from reduces the distance to if for any (we do not assume that ). In other words, reduces the distance to if
For example, the path from
reduces the distance to Note that, if and is a path from reducing the distance to we have for small enough
Corollary 4.16.
Let and If such that and there is a path from reducing the distance to then In particular, if for any distinct there is either a path from reducing the distance to or a path from reducing the distance to then
Example 4.17.
Consider the standard circle By Corollary 4.16 for any we have
Example 4.18.
Let be an arc lying in one open semicircle.
In other words the anticlockwise directed distance from to is strictly less than Then by Corollary 4.16 we obtain for any
Example 4.19.
Now assume that is an arc that does not lie in one semicircle.
So the anticlockwise directed distance from to is strictly greater than Moreover, let us assume that the points are close enough Then we claim that for
(in this case has finite rank, while has infinite rank). Let us prove this.
If then for any such that there is either a path from reducing the distance to or a path from reducing the distance to So
Now assume that Then all pairs except and lie in because there is a path from reducing the distance to We claim that For this we write everything in coordinates in a symmetric form and If using that and one can show that is in the Quadrant two and is in the Quadrant one.
Then there is no a point such that and Hence This also implies that is not equivalent to Therefore